# ExMAG: Learning of Maximally Ancestral Graphs

### v2

@@ 4.2 MILP Formulation
$l_{q}$ $=\sum_{i=1}^{n}\sum_{j=1}^{d}\left[X_{i,j}-\sum_{k=0;k\neq j}^{d}X_{i,k}w_{k,j}\right]^{q}$ (13)
$\quad+\lambda\sum_{j=0}^{d}\sum_{k=0}^{d}(e_{j,k}+b_{j,k}),$
@@
  • $b_{j,k}$ is the binary decision variable indicating the presence of a bidirected edge between $j$ and $k$;
  • $\lambda$ is a regularization parameter controlling the trade-off between the model fit and the edge penalty.
  • The exponent $q\in\mathbb{N}$ can take values $q=1$ or $q=2$. While setting $q=1$ leads to a mixed-integer linear program (MILP), which case would not correspond to the maximum likelihood estimator. If we choose $q=2$, resulting in a mixed-integer quadratic problem(MIQP).
As in the ExDAG model, the weights are bounded by introducing a large constant $c$ that ensures the weights are appropriately constrained when there is an edge between the variables. The bounding avoids bilinear terms in the cost function in (13). As a matrix form, the weight constraints can be restrained as:
@@
The (4.2) means that there cannot be a directed as well as a bidirected edge between the same two vertices. Additionally, we enforce that the bidirected matrix is symmetric, i.e.,
$B=B^{T}.$ (14)
If $f_{j,k}=1$, then we know there is no direct causal relationship between $j$ and $k$, and therefore, $e_{j,k}=0$. Formally, this translates to condition
$F+E\leq\mathbf{1}.$ (15)
Inversely, $f_{j,k}=0$ implies a directed edge rather than a bidirected edge between $j$ and $k$. Therefore,
@@
Lastly, we must enforce conditions for directed or almost directed cycles and inducing paths. Those conditions are enforced lazily using a separation routine explained later. Directed cycles are enforced in a way adopted from [46]. Therefore, they are left out of this paper. Suppose an almost directed cycle is formed by edges in set $E$ and a bidirected edge $(u,v)$. Then, this almost directed cycle is forbidden by constraint
$b_{u,v}+\sum_{(j,k)\in E}e_{j,k}\leq|E|.$
Similarly, if there is an inducing path formed by path $P$ that contains bidirected edges, and set $E$ contains all directed edges that participate in the ancestor relationship (including multiple paths) between the inner points of the path and the terminals of $P$, this inducing path is forbidden by
$\sum_{(j,k)\in P}b_{j,k}+\sum_{(j,k)\in E}e_{j,k}\leq|E|+|P|-1.$
Note that the second condition does not necessarily eliminate the inducing path, as the optimizer might forbid one of the directed edges in $E$ without influencing the ancestor relationship. This results in the path $P$ being found in the next iteration, with a smaller set of directed edges, and the process is repeated.
@@ 5 Separation Routine for the Maximal Ancestral Graphs
The main contribution of this section is the separation routine that allows us to identify whenever a graph is not an instance of a maximal ancestral graph. To do so, we need to identify directed cycles, almost directed cycles, and inducing paths. The directed cycles can be found in $\mathcal{O}(d^{2})$ using depth-first-search (DFS); such an approach can be found in [46]. For both inducing paths and almost directed cycles, we will use a distance matrix constructed on the graph of directed edges. Such distance matrix can be obtained using the Floyd-Warshall algorithm [20].
Once having the distance matrix, to check for almost directed cycles, we can iterate over all bidirected edges and test whether the distance between the endpoints is finite, i.e., we have a directed path connected by a bidirected edge. See Algorithm 1 for details.
@@
     for all $j\in{1,2,\ldots,d}$ do
         for all $k\in{1,2,\ldots,d}$ do
              if $b_{j,k}==1\And d_{j,k}<\infty$ then
                  $E=$ Trace-Distance-Matrix($d$,$e$,$j$,$k$)
                  Found cycle formed by edges $E$ and $j\leftrightarrow k$
@@
Algorithm 1 Function to identify almost directed cycles.
In the case of inducing paths, the problem is more complicated. We use DFS started from each of the vertices. Once started from vertex $x$, this DFS routine checks for all possible inducing paths that terminate in $x$. For efficiency, a set of all possible endpoints of the path is held. Once this set is empty, the DFS search is terminated, and no further exploration is performed. The set is updated using the distance matrix calculated from the graph of the directed edges. If we consider a vertex $y$, $x$ must be either its ancestor (meaning that possible endpoints for $y$ remain unchanged) or it is among the points that are reachable from $y$ (meaning that the possible endpoints for $y$ are replaced with their intersection with the set of all points reachable from $y$). See Algorithm 2 for details.
@@
     for all $j\in{1,2,\ldots,d}$ do
         Inducing-Paths-DFS($d$, $e$, $b$, $j$, $j$, $\{1,2,\ldots d\}$, $[j]$)
@@
function Inducing-Paths-DFS($d$, $e$, $b$, start, $u$, possible endpoints, path)
@@
     if Len(path) $>2\And u$ in possible endpoints  then
         Found-Inducing-Path($d$,$e$, path)
@@
         if $d_{v,start}<\infty$ then v-endpoints = possible endpoints
         elsev-endpoints = possible endpoints $\cap\{x\mid d_{v,x}<\infty\}$
         end if
         Inducing-Paths-DFS($d$,$e$,$b$,start,$v$,v-endpoints,path$+v$)
@@
function Found-Inducing-Path($d$,$e$, $P$)
[…]
    for all vertices $j\in$ $P$ and $j\not\in\{P_{0},P_{|P|}\}$ do
        $E=E\cup$ Trace-Distance-Matrix($d$,$e$,$j$,$P_{0}$)
        $E=E\cup$ Trace-Distance-Matrix($d$,$e$,$j$,$P_{|P|}$)
    end for
    Found inducing path formed by path $P$ and directed edges $E$
end function
@@
Once having the bidirected edges in the inducing paths and almost directed cycles, we need to trace back the Floyd-Warshall distance matrix to find all edges that form the cycle or the ancestor relationship. This is done using calls to the function in Algorithm 3.
If directed cycles, almost directed cycles, and inducing paths are found, the algorithm applies lazy constraints in 4.2 and in 4.2. Otherwise, we know that the program converged to the optimum, and we have a maximal ancestral graph, which minimizes (13).
Output: edges on any path from $j$ to $k$
function Trace-Distance-Matrix($d$, $e$, $j$, $k$)
     if $d_{j,k}==\infty$ then return
     end if
     visited = $\{(j,k)\}$
     stack = stack with $(j,k)$
[…]
     while stack is not empty do
         $u,v\leftarrow$ Pop(stack)
         for all $w\in{1,2,\ldots,d}$ s t. $d_{u,w}+d_{w,v}<\infty$ do
              visited $\leftarrow$ visited $\cup\{(u,w),(w,v)\}$
              if $e_{u,w}$ then edges $\leftarrow$ edges $\cup\{(u,w)\}$
              else if $e_{w,v}$ then edges $\leftarrow$ edges $\cup\{(w,v)\}$
              end if
              add $\{(u,w),(w,v)\}$ to stack
         end for
     end while
end function
Algorithm 3 Functions that help in the separation routine.
@@ 6 Experimental Evaluation
Figure 2: SHD values (in the vertical axis) for different settings of $d$ (in the horizontal axis) and $n$ (horizontal choice of the graph). The plots in the vertical dimension differ according to the dataset used. The channel shows the variance of the results for different choices of seed. See supplementary materials for results on more datasets.
6.1 Used Datasets
@@
We test the ExMAG algorithm on both synthetic and real-world datasets. The first synthetic dataset is based on the Erdös-Rényi model (ER) [19], in which the ground truth graph is randomly selected from all graphs with $d$ vertices and $m$ edges (parameter of the experiment, for example, dataset ER-5 contains $5$ edges per variable, i.e., $m=5\cdot d$). The weights of the graph are randomly sampled form the set $(-2.0,-0.5)\cup(0.5,2.0)$.
@@
Once the ground truth model is created, then the training data are generated using the structural model equation 8,  9 ($H_{i}$ set to 00). Then, $20\,\%$ of variables are treated as latent variables and hidden from the training data. The respective columns and rows from the ground truth weight matrix $W$ have also been removed. Finally, $20\,\%$ of edges between variables not connected by an edge in the ground truth data are marked in $F$.
@@
The second dataset uses randomly generated bow-free (BF) graphs, a subset of all possible MAGs. A bow-free graph is defined as a graph that does not contain a pair of vertices $i,j$, such that $i\rightarrow j$ and at the same time $i\leftrightarrow j$. The BF graph generation process has two parameters. First, the probability of a directed edge. Second, the probability of bidirected edge. The generation process is as follows. First, a bow-free graph with the given edge probabilities is generated randomly. The weights of the sampled graphs are randomly sampled from the set $(-2.0,-0.5)\cup(0.5,2.0)$. The adjacency matrix of the directed edges defines the weights of the structural equation model. Then the data samples are generated using the structural equation, where the noise is sampled from multivariete Gaussian distribution with covariance matrix equals to the adjacency matrix of bidirected edge generated in the previous step.
@@
The third dataset uses real-world data from the financial sector. Paper [1] works with systemic credit risk, one of the most important concerns within the financial system, using dynamic Bayesian networks. The data show that transport and manufacturing companies are likely to transfer risk to other sectors, while banks and the energy sector are likely to be influenced by the risks from other sectors. The data from [1] contains a 10-time series capturing the spreads of 10 European credit default swaps (CDS), and further six time series are added from [45].
@@
6.2 Evaluation Criteria
Suppose that a tested algorithm produced weight matrix $\hat{W}$. Such a matrix can contain nearly zero weights. For such reasons, thresholding is done, keeping only edges with weight greater or equal to $\delta$. In cases when the ground-truth weight matrix $W$ is known, the best solution (in terms of structural Hamming distance, see below) is kept over those defined by different threshold $\delta$ values.
In the evaluation, we use the structural Hamming distance (SHD). This distance is the sum of contributions over all pairs of variables in the graph. For two variables $i,j$, let $GT\in\{\rightarrow,\leftarrow,\leftrightarrow,\emptyset\}$ be the edge type in the ground truth graph and $PR\in\{\rightarrow,\leftarrow,\leftrightarrow,\emptyset\}$ be edge type in the predicted graph. Then the contribution of $i,j$ pair to SHD is
$r_{ij}=\begin{cases}0,&\mbox{if }GT=PR,\\ 0.5,&\mbox{if }GT\neq PR\mbox{ and }GT\neq\emptyset\mbox{ and }PR\neq\emptyset,\\ 1,&\mbox{otherwise.}\end{cases}$
Other measured criteria include runtime and F1-score, i.e., the harmonic mean of precision and recall.
6.3 Experiment Setting
In the experiments, we show the results of ExMAG. In the case of synthetic datasets, we generated random graphs with $d\in\{5,10,15,20,25\}$. The number of samples was in $n\in\{20,50,100,500,1000\}$, and ground-truth graph edge ratio was in $\{2,3,4,5,6,7,8,9,10,15,20\}$ percent. Both tested algorithms were run $10$ times, each time with a different seed value and the results were then averaged.
6.4 Experimental Results
The SHD results is shown in Fig. 2 and in the supplementary materials on more datasets. The plots show a comparison of SHD values for ExMAG on both synthetic datasets. As can be seen, the structural Hamming distance grows with the number of variables and decreases with the number of samples. For results with the F1-score, visit the supplementary materials. The results show that the F1 score is close to $1$, with the score increasing with the higher number of samples and decreasing with the higher number of variables.
The results on the real-world dataset can be seen in 3. Contrary to the original expectations, the highest risk importer is company 2H66B7, which stands for Lufthansa. The second highest risk importer is 2H6677, i.e., the Deutsche Bank, which is an expected result.
7 Conclusion
Learning of Dynamic Bayesian networks has received considerable attention as a means of causal learning. With a few exceptions, the research has not considered confounding explicitly. Our method, ExMAG, estimates a maximally ancestral graph, capturing confounding and causal relationships using bidirected and directed edges of a mixed graph. The method provides state-of-the-art statistical performance.
The predictive power of forecasting with variants of dynamical Bayesian networks with confounding considerations is an important direction for further work. While it seems clear that (1) even in dynamical Bayesian networks, marginalization is hard, and thus the computational complexity may be high, (2) the statistical performance should improve by considering the confounding.

### v3

@@ 3.2 MILP Formulation
$l_{q}$ $=\sum_{i=1}^{n}\sum_{j=1}^{d}\left[X_{i,j}-\sum_{k=0;k\neq j}^{d}X_{i,k}w_{k,j}\right]^{q}+\lambda\sum_{j=0}^{d}\sum_{k=0}^{d}(e_{j,k}+b_{j,k}),$ (4)
@@
  • $b_{j,k}$ is the binary decision variable indicating a bidirected edge between $j$ and $k$;
  • $\lambda$ is a regularization parameter controlling the model fit and the edge penalty trade-off.
  • The exponent $q\in\mathbb{N}$ can take values $q=1$ or $q=2$.
As in the ExDAG [48] model, the weights are bounded by introducing a large constant $c$.The bounding avoids bilinear terms in the cost function in (4) and takes the following form:
@@
The 3.2 means that there cannot be a directed as well as a bidirected edge between the same two vertices. Additionally, we enforce that the bidirected matrix is symmetric by (5). If $f_{j,k}=1$, then we know there is no direct causal relationship between $j$ and $k$, and therefore, $e_{j,k}=0$. This is formally enforced by (6) Inversely, $f_{j,k}=0$ implies a directed edge rather than a bidirected edge between $j$ and $k$ in (7).
$B$ $=B^{T},$ (5)
$F+E$ $\leq\mathbf{1},$ (6)
@@
Lastly, we must enforce conditions for directed or almost directed cycles and inducing paths. Those conditions are enforced lazily using a separation routine explained later. Directed cycles are enforced in a way adopted from [49]. Therefore, they are left out of this paper. An almost directed cycle formed by edges in set $E^{\prime}$ and a bidirected edge $(u,v)$ is forbidden by the constraint
$b_{u,v}+\sum_{(j,k)\in E^{\prime}}e_{j,k}\leq|E^{\prime}|.$
Similarly, if there is an inducing path formed by path $P$ that contains bidirected edges, and set $E^{\prime}$ contains all directed edges that participate in the ancestor relationship (including multiple paths) between the inner points of the path and the terminals of $P$, this inducing path is forbidden by
$\sum_{(j,k)\in P}b_{j,k}+\sum_{(j,k)\in E^{\prime}}e_{j,k}\leq|E^{\prime}|+|P|-1.$
Note that the second condition does not necessarily eliminate the inducing path, as the optimizer might forbid one of the edges in $E^{\prime}$ without influencing the ancestor relationship. This results in path $P$ being found in the next iteration, with a smaller set of directed edges, and the process is repeated.
@@ 4 Separation Routine for the Maximal Ancestral Graphs
The main contribution of this section is the separation routine that identifies whenever a graph is an instance of a maximal ancestral graph. To do so, we need to identify directed cycles, almost directed cycles, and inducing paths. The presence of directed cycles can be detected in $\mathcal{O}(d^{2})$ using depth-first-search (DFS); such an approach can be found in [49]. For both inducing paths and almost directed cycles, we will use the distance matrix $D$ constructed on the graph of directed edges $E$. This distance matrix can be obtained, for example, using the Floyd-Warshall algorithm [22].
Having the distance matrix, to check for almost directed cycles, we can iterate over all bidirected edges and test whether the distance between the endpoints using $E$ is finite, i.e., we have a directed path connected by a bidirected edge. See Algorithm 1 for details.
@@
     for all $(j,k)\in\{1,2,\ldots,d\}\times\{1,2,\ldots,d\}$ do
         if $j\neq k\And b_{j,k}==1\And D_{j,k}<\infty$ then
              $E^{\prime}=$ Trace-Distance-Matrix($D$,$E$,$j$,$k$) $\triangleright$ Finds all edges on any $j$ to $k$ path, see Supl.
              Found cycle formed by edges $E^{\prime}$ and $j\leftrightarrow k$
@@
Algorithm 1 Function that identifies almost directed cycles.
In the case of inducing paths, we use a DFS starting from each vertex. Once started from vertex $s$, the DFS routine checks for all possible inducing paths that terminate in $s$. For efficiency, a set of all possible endpoints of the path is held. Once this set is empty, the DFS search is terminated, and no further exploration is performed. The set is updated using the distance matrix calculated on the directed edges. If we consider a vertex $v$, $s$ must be either its ancestor (meaning that possible endpoints for $v$ remain unchanged) or the second inducing path endpoint is among the points that are reachable from $v$ (meaning that the possible endpoints for $v$ are replaced with their intersection with the set of all points reachable from $v$). See Algorithm 2 for details.
@@
     for all $s\in{1,2,\ldots,d}$ do Inducing-Paths-DFS($D$, $E$, $B$, $s$, $s$, $\{1,2,\ldots d\}$, $[s]$)
@@
function Inducing-Paths-DFS($D$, $E$, $B$, $s$, $u$, possible endpoints, path)
@@
     if Len(path) $>2\And u$ in possible endpoints  then Found-Inducing-Path($D$,$E$, path)
         $\triangleright$ Edges participating in ancestor relation are recovered, see Suppl. mat.
@@
         v-endpoints $\leftarrow$ possible endpoints if $D_{v,s}<\infty$ else possible endpoints $\cap\{x\mid D_{v,x}<\infty\}$
         Inducing-Paths-DFS($D$,$E$,$B$,$s$,$v$,v-endpoints,path$+v$)
@@
Once having the bidirected edges in the inducing paths and almost directed cycles, we need to trace back the Floyd-Warshall distance matrix to find all directed edges that form the cycle or the ancestor relationship. This is done using calls to the function Trace-Distance-Matrix, which can be found in the Supplementary materials.
If directed cycles, almost directed cycles, and inducing paths are found, the algorithm applies lazy constraints in 3.2 and in 3.2. Note that removing one directed edge between two vertices where multiple paths exist is not a necessary condition for the graph to become MAG; however, this procedure can be repeated iteratively. If no inducing paths or almost directed cycles are found, we know that the program converged to the optimum, and we have a maximal ancestral graph, which minimizes (4).
@@ 5 Experimental Evaluation
Figure 2: SHD values (in the vertical axis) for different settings of $d$ (in the horizontal axis) and $n$ (horizontal choice of the graph). The plots in the vertical dimension differ according to the dataset used. Standard deviations are depicted as the blurred regions, and dashed lines are the maximum values. See supplementary materials for results on more datasets and error information.
Figure 3: SHD values on 3BF datasets.
Figure 4: Run times of the compared algorithms in seconds.
Used Datasets
@@
We tested the ExMAG algorithm on both synthetic and real-world datasets. The first synthetic data set is based on the Erdös-Rényi model (ER) [21], in which the ground truth graph is randomly selected from all graphs with $d$ vertices and $m$ edges (parameter of the experiment, for example, dataset ER-2 contains $2$ edges per variable, that is, $m=2\cdot d$). The weights of the graph are randomly sampled from the set $(-2.0,-0.5)\cup(0.5,2.0)$.
@@
Once the ground truth model is created, the training data are generated using the structural model equation (2),  (3) ($H_{i}$ set to 00). Then, $20\,\%$ of variables are treated as latent variables and hidden from the training data. The respective columns and rows from the ground truth weight matrix $W$ have also been removed. Finally, $20\,\%$ of edges between variables not connected by an edge in the ground truth data are marked in $F$.
@@
The second dataset uses randomly generated bow-free (BF) graphs, a subset of all possible MAGs. A bow-free graph is a graph such that for no pair of vertices $i,j$, $i\rightarrow j$ and at the same time $i\leftrightarrow j$. The BF graph generation process has two parameters: the probability of a directed edge and the probability of a bidirected edge. The generation process is as follows. First, a bow-free graph with the given edge probabilities is generated randomly. Then the weights of the sampled graphs are randomly sampled from the set $(-2.0,-0.5)\cup(0.5,2.0)$.
The third synthetic dataset, 3BF, is a modified version of BF. We generate the ground truth graph in the same way as in the BF dataset and then modify it. Specifically, we identify each vertex with a degree greater than three, and randomly remove edges until the vertex has a degree of at most three.
The adjacency matrix of the directed edges defines the weights of the structural equation model. Then the data samples are generated using the structural equation, where the noise is sampled from a multivariate Gaussian distribution with a covariance matrix equal to the adjacency matrix of bidirected edges generated in the previous step.
@@
The fourth dataset uses real-world data from the financial sector. Paper [1] works with systemic credit risk, one of the most important concerns within the financial system, using dynamic Bayesian networks. The data show that transport and manufacturing companies are likely to transfer risk to other sectors, while banks and the energy sector are likely to be influenced by the risks from other sectors. The data from [1] contains a 10-time series capturing the spreads of 10 European credit default swaps (CDS), and further six time series are added from [48].
Evaluation Criteria
Suppose that a tested algorithm produced weight matrix $\hat{W}$. Such a matrix can contain nearly zero weights. For such reasons, thresholding is done, keeping only edges with weight greater than or equal to $\delta$. In cases when the ground-truth weight matrix $W$ is known, the best solution (in terms of structural Hamming distance, see below) is kept over those defined by different threshold $\delta$ values. In the evaluation, we use the structural Hamming distance (SHD). This distance is the sum of contributions over all pairs of variables in the graph. For two variables $i,j$, let $GT\in\{\rightarrow,\leftarrow,\leftrightarrow,\emptyset\}$ be the edge type in the ground truth graph and $PR\in\{\rightarrow,\leftarrow,\leftrightarrow,\emptyset\}$ be edge type in the predicted graph. Then the contribution of $i,j$ pair to SHD is $r_{ij}=0$ if $GT=PR$, $0.5$ if $GT\neq PR\wedge GT\neq\emptyset\wedge PR\neq\emptyset$, and $1$ otherwise. Other measured criteria include runtime and F1-score, i.e., the harmonic mean of precision and recall.
Experiment Setting
In the experiments, we show the results of ExMAG. In the case of synthetic datasets, we generated random graphs with the number of vertices $d\in\{5,10,15,20,25\}$. The number of samples was $n\in\{20,50,100,500,1000\}$, and for the ground-truth graph, the edge-to-vertex ratio was in $\{2,3,4,5,6,7,8,9,10,15,20\}$. All tested algorithms were run $10$ times, each time on synthetic data generated using a random generator initialized with a different seed. The results were then averaged. We compared our method with the FCI algorithm [51], ip4ancadmg [11], and ipbm [17]. We set regularization coefficient $\lambda$ to $1.0$. We ran experiments on a computing cluster with AMD EPYC 7543 cpus and each job had allocated two cores and 64GB RAM. Time limit was 900 seconds for ExMAG and 1800 seconds for other methods. The total cpu time needed for experiments in this paper was around one month.
@@
Experimental Results
The SHD results are shown in Figures 2 and 3 and in the supplementary materials for additional datasets. The plots show a comparison of SHD values for ExMAG on the synthetic datasets. As can be seen, the structural Hamming distance grows with the number of variables and decreases with the number of samples.
As we can see on Figure 2 and 3, ExMAG performs better than FCI on all scenarios. Since both ipbm and ip4ancadmg have a preprocessing step that depends exponentially on the maximum in-degree of the underlying ground truth graph, we tested these two algorithms only on the 3BF datasets, where the in-degree is bounded by three. We can see in Figure 3, that ExMAG also performs better than ipbm and ip4ancadmg. The run times of the evaluated algorithms are shown in Figure 4. For additional results (incl. the F1-score), please see the supplementary materials.
The results on the real-world dataset can be seen in Figure 5. Contrary to the original expectations, the highest risk importer is company 2H66B7, which stands for Lufthansa. The second highest risk importer is 2H6677, i.e., the Deutsche Bank, which is an expected result.
6 Conclusion and Limitations
Learning of dynamic Bayesian networks has received considerable attention as a means of causal learning. With a few exceptions, the research has not considered confounding explicitly. Our method, ExMAG, estimates a maximally ancestral graph, capturing confounding and causal relationships using bidirected and directed edges in a mixed graph. The method provides state-of-the-art statistical performance.
As with many other methods for causal learning, the scalability of the method may leave space for improvement. Although the branch-and-cut algorithm runs in time that is exponential in the number of time series in the worst case, Figure 4 illustrates that our run time is lower than those of ip4ancadmg [11] and ipbm [17], while improving the SHD of both recent competitors at the same time (cf. Figure 3). One could improve upon the run-time further by introducing additional cutting planes and more elaborate data structures for the separation of 3.2 and 3.2, perhaps drawing inspiration from solvers [15, e.g.] for the travelling salesman problem.
In terms of future work, exploring the predictive power of forecasting using variants of dynamical Bayesian networks with confounding considerations seems prominent. Although it seems clear that marginalization is hard even in dynamical Bayesian networks, and thus the computational complexity may be high, but statistical performance is likely to improve, when confounding is considered.